OpenStudy (anonymous):

Solve. (x – 5)^2 = 4.

OpenStudy (anonymous):

x^2 -10x +25 = 4 x^2 -10x + 21=0 (x-7)(x-3) = 0 x=7,3

OpenStudy (anonymous):

it's easier to square root both sides since the 4 is a perfect square. This will give you (x-5)=2/-2 then add 5 to both sides x=7/3